OpenStudy (anonymous):

Perform the indicated operation g(a) = -3a + 1 h(a) = 3a3 + 4a @iGreen

OpenStudy (igreen):

You need the solution to those two equations?

OpenStudy (anonymous):

its said find g(n) h(n)

OpenStudy (igreen):

I'm not sure..sorry.

jigglypuff314 (jigglypuff314):

g(n) h(n) = ? I might be wrong, but my best guess would be to multiply them together and foil? (-3x+1) (3a^3+4a) do you know how to foil that @morrisd7

OpenStudy (anonymous):

no i dont know foil

jigglypuff314 (jigglypuff314):

foil is the distributive property that includes polynomials FOIL stands for First, Outer, Inner, Last (aka the order you multiply) so (a+b)(c+d) = ac + ad + bc + bd so then (-3a+1) (3a^3+4a) = (-3a)(3a^3) + (-3a)(4a) + (1)(3a^3) + (1)(4a) can you simplify that now?

OpenStudy (anonymous):

no

jigglypuff314 (jigglypuff314):

hmmm if you have not learned this then are you sure it said to find "g(n) h(n)" or was it "g(n)+h(n)" something like that?

OpenStudy (anonymous):

yea

jigglypuff314 (jigglypuff314):

(-3a)(3a^3) + (-3a)(4a) + (1)(3a^3) + (1)(4a) becomes -9a^4 - 12a^2 + 3a^3 + 4a